40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Mar 18 10:46:51 2010, Koji, Update, ABSL, PLL reconstructed 
    Reply  Thu Mar 18 12:27:56 2010, Alberto, Update, ABSL, PLL reconstructed 
       Reply  Thu Mar 18 13:40:35 2010, Koji, Update, ABSL, PLL reconstructed 
          Reply  Thu Mar 18 21:42:26 2010, Koji, Update, ABSL, PLL reconstructed 
             Reply  Mon Mar 22 22:11:26 2010, Mott, Update, ABSL, PLL reconstructed IMG_0150.JPG
                Reply  Mon Mar 22 23:37:32 2010, Mott, Update, ABSL, PLL reconstructed 2010-03-22_23.14.00.jpg2010-03-22_23.24.50.jpg
                   Reply  Tue Mar 23 18:44:46 2010, Mott, Update, ABSL, PLL reconstructed NPRO_PLL_freqresp.png
                      Reply  Wed Mar 24 03:58:18 2010, kiwamu, matt, koji, Update, Green Locking, leave PLL locked DSC_1402.JPG
                         Reply  Wed Mar 24 04:22:51 2010, kiwamu, matt, Update, Green Locking, two NPRO PLL  
                            Reply  Wed Mar 24 14:52:02 2010, Alberto, Update, Green Locking, two NPRO PLL  
                         Reply  Wed Mar 24 15:59:59 2010, kiwamu, matt, Update, Green Locking, leave PLL locked 
                            Reply  Thu Mar 25 17:29:48 2010, kiwamu, mott, Update, Green Locking, PLL two NPROs 
Message ID: 2706     Entry time: Wed Mar 24 03:58:18 2010     In reply to: 2703     Reply to this: 2707   2712
Author: kiwamu, matt, koji 
Type: Update 
Category: Green Locking 
Subject: leave PLL locked 

We are leaving the PLL as it is locked in order to see the long term stability. And we will check the results in early morning of tomorrow.

DO NOT disturb our PLL !!

  


(what we did)

After Mott left, Matt and I started to put feedback signals to the temperature control of NPRO.

During doing some trials Matt found that NPRO temperature control input has an input resistance of 10kOhm.

Then we put a flat filter ( just a voltage divider made by a resistor of ~300kOhm and the input impedance ) with a gain of 0.03 for the temperature control to inject a relatively small signal, and we could get the lock with the pzt feedback and it.

In addition, to obtain more stable lock we then also tried to put an integration filter which can have more gain below 0.5Hz.

After some iterations we finally made a right filter which is shown in the attached picture and succeeded in obtaining stable lock.

 

 

 

Attachment 1: DSC_1402.JPG  1.125 MB  Uploaded Wed Mar 24 05:00:01 2010  | Hide | Hide all
DSC_1402.JPG
ELOG V3.1.3-